Equivalent Units of Production and Related Costs
The charges to Work in Process—Assembly Department for a period, together with information concerning production, are as follows. All direct materials are placed in process at the beginning of production.

Work in Process—Assembly Department
Bal., 9,000 units, 40% completed
72,360
To Finished Goods, 41,500 units
?
Direct materials, 40,000 units @ $6.80
272,000


Direct labor
80,000


Factory overhead
40,450


Bal., ? units, 30% completed
?


Determine the following:
a. The number of units in work in process inventory at the end of the period.
units

Feedback
b. Equivalent units of production for direct materials and conversion. If an amount is zero or a blank, enter in "0".
Work in Process-Assembly Department
Equivalent Units of Production for Direct Materials and Conversion Costs


Whole
Units
Equivalent Units
Direct Materials
Equivalent Units
Conversion
Inventory in process, beginning

Started and completed

Transferred to finished goods

Inventory in process, ending

Total units

c. Costs per equivalent unit for direct materials and conversion.

Costs Per Equivalent Unit
Direct Materials
$
Conversion
$
d. Cost of the units started and completed during the period.
$
Answer #1

a) Ending work in process = 9000+40000-41500 = 7500

b) Equivalent unit

Whole unit Direct material equivalent unit Conversion cost equivalent unit
Inventory in process, beginning 9000 0 9000*60% = 5400
Unit started and completed 32500 32500 32500
Unit transferred out 41500 32500 37900
Inventory in process, ending 7500 7500 7500*30% = 2250
Total 49000 40000 40150

c) Cost per equivalent unit of material = 272000/40000 = 6.80

Cost per equivalent unit of conversion = (80000+40450)/40150 = 3

d) Cost of unit started and completed = 9.80*32500 = 318500